A student with a mass of 50.1 kg jumps straight up in the air by using her legs to apply an average force of 847 N to the ground for 0.389 s. Assume that the initial momentum of the student and the Earth are zero. (The mass of the Earth is 5.98·1024 kg.)

  1. What is the momentum of the student immediately after this impulse? (Convention: upwards is positive and downwards is negative.)
  2. What is the momentum of the Earth after this impulse? (Convention: upwards is positive and downwards is negative.)
  3. What is the speed of the Earth right after the impulse?
  4. What fraction of the total kinetic energy that the student produces with her legs goes to the Earth?
  5. Using conservation of energy, how high does the student jump?
Answer #1

a)

Psi = initial momentum of the student = 0 kgm/s

Psf = final momentum of the student = ?

F = average force = 847 N

t = time = 0.389 s

Using Impulse- change in momentum equation

F t = Psf - Psi

847 x 0.389 = Psf - 0

Psf = 329.5 kgm/s

b)

Pei = initial momentum of earth = 0 kgm/s

Pef = final momentum of earth

Using conservation of momentum

Pef + Psf = Pei + Psi

Pef + 329.5 = 0 + 0

Pef = - 329.5 kgm/s

c)

ve = speed of earth = ?

me = mass of earth = 5.98 x 1024 kg

Final momentum of earth is given as

Pef = me ve

329.5 = (5.98 x 1024) ve

ve = 5.51 x 10-23 m/s

d)

Final momentum of student is given as

Psf = ms vs

329.5 = (50.1) vs

vs = 6.6 m/s

vs = 5.51 x 10-23 m/s

Total kinetic energy is given as

KEt = (0.5) me v2e + (0.5) ms v2s

kinetic energy of earth is given as

KEe = kinetic energy of earth = (0.5) me v2e

Fraction of total kinetic energy that goes to earth is given as

KEe /KEs = (0.5) me v2e /((0.5) ms v2s ) = (5.98 x 1024) ( 5.51 x 10-23)2 /((50.1) (6.6)2) = 8.32 x 10-24

e)

h = height gained by student

Height gained by student is given as

h = v2s /(2g)

h = (6.6)2/(2 x 9.8)

h = 2.22 m